OpenStudy (anonymous):

how long would it take tim and jim to mow a lawn together if tim usually mows it in 6 hrs and jim in 7?

OpenStudy (anonymous):

tim can do 1/6 of the job in 1 hours jim can do 1/7 of the jobs in 1 hours (1/6+1/7)h=1 find h, hours

OpenStudy (anonymous):

\[\frac{7\times 6}{7+6}\]

OpenStudy (anonymous):

if you solve what iman said for h that is what you will get. his equation was correct.

OpenStudy (anonymous):

but if you see this problem a lot (two people only) you can go right to the answer.

OpenStudy (anonymous):

\[\frac{1}{6}+\frac{1}{7}=\frac{1}{x}\]\[x=\frac{42}{13}\text{hours} \]

OpenStudy (anonymous):

3.23 hrs?

OpenStudy (anonymous):

3.23077

OpenStudy (anonymous):

Probably About 3 Hours And 30 Minutes.?

OpenStudy (anonymous):

3hrs 13minutes 50.772 seconds